CUET Mathematics: AlgebraProbability. Free, no login required.

Q1:

2025: 3 June Shift 2

Probability

Medium

Applied

The curve y=f(x)y = f(x) is normal probability curve, then which of the following statements are correct?

(A) mean, median and mode of the distribution coincide.

(B) the area bounded by the curve y=f(x)y = f(x) and xx-axis is one unit.

(C) The curve is symmetrical about the line x=μx = \mu, where μ\mu is the mean.

(D) yy-axis is an asymptote to the curve.

Choose the correct answer from the options given below:

Answer options
Option 1
Correct Answer
Explanation for 2025: 3 June Shift 2 MAT question 1

Q2:

2025: 3 June Shift 1

Probability

Easy

Common

A random variable X has the following probability distribution:

X2345
P(X)5/k7/k9/k11/k

Then the value of k4\frac{k}{4} is:

Answer options
Option 1
Correct Answer
Explanation for 2025: 3 June Shift 1 MAT question 2

Q3:

2025: 3 June Shift 1

Probability

Easy

Core

60% members of a committee favour a certain proposal and 40% members oppose the proposal. A member is selected and let the random variable X = 0 if he opposes and X = 1 if he is in favour. Then the variance of the random variable X is

Answer options
Option 3
Correct Answer
Explanation for 2025: 3 June Shift 1 MAT question 3

Q4:

2025: 3 June Shift 1

Probability

Medium

Core

If A and B are independent events, then which of the following is not true?

Answer options
Option 4
Correct Answer
Explanation for 2025: 3 June Shift 1 MAT question 4

Q5:

2025: 3 June Shift 1

Probability

Medium

Core

The probability that a leap year selected at random will have 53 Mondays is

Answer options
Option 2
Correct Answer
Explanation for 2025: 3 June Shift 1 MAT question 5

Q6:

2025: 3 June Shift 1

Probability

Medium

Core

Two persons A and B throw a die alternately till one of them gets a 'three' and wins the game. The probability of A's winning if A starts first is

Answer options
Option 1
Correct Answer
Explanation for 2025: 3 June Shift 1 MAT question 6

Q7:

2025: 3 June Shift 1

Probability

Medium

Applied

If the difference between mean and variance of a Binomial distribution is 1 and the difference of their squares is 5, then the probability of success is

Answer options
Option 2
Correct Answer
Explanation for 2025: 3 June Shift 1 MAT question 7

Q8:

2025: 3 June Shift 1

Probability

Medium

Applied

Let X denote the number of hours a student studies on a selected day. The probability distribution of X is given by (where k is some unknown constant)

P(X=xi)={0.5,if xi=0,kxi,if xi=1,k(4−xi),if xi=2 or 3,0,otherwise.P(X = x_i) = \begin{cases} 0.5, & \text{if } x_i = 0, \\ kx_i, & \text{if } x_i = 1, \\ k(4 - x_i), & \text{if } x_i = 2 \text{ or } 3, \\ 0, & \text{otherwise}. \end{cases}

Then the value of k is

Answer options
Option 3
Correct Answer
Explanation for 2025: 3 June Shift 1 MAT question 8

Q9:

2025: 3 June Shift 1

Probability

Medium

Applied

Two percent of the bolts manufactured in a factory are found to be defective. Using the Poisson distribution, the probability that in a sample of 100 bolts chosen at random, exactly two will be defective, is: [Given e−2=0.135e^{-2}=0.135]

Answer options
Option 4
Correct Answer
Explanation for 2025: 3 June Shift 1 MAT question 9

Q10:

2025: 3 June Shift 1

Probability

Medium

Applied

Which of the following statements are correct?

(A) The mean and variance of the Poisson distribution are equal.

(B) The mean and variance of a Binomial distribution are equal.

(C) An unbiased die is thrown again and again until two sixes are obtained, then the probability of obtaining the second six in the 3rd throw is 5108\frac{5}{108}.

(D) If the variance of a Poisson distribution is 2, then P(X = 2) = 2e−22e^{-2}

Choose the correct answer from the options given below:

Answer options

Q11:

2025: 2 June Shift 1

Probability

Medium

Common

The probability distribution of a random variable xx is, P(x)=k2x,x=0,1,2,3P(x) = \frac{k}{2^x}, x = 0, 1, 2, 3. Then

Match List-I with List-II

List-IList-II
(A) kk(I) 215\frac{2}{15}
(B) P(x=1)P(x = 1)(II) 15\frac{1}{5}
(C) P(1<x<3)P(1 < x < 3)(III) 815\frac{8}{15}
(D) P(x≥2)P(x \geq 2)(IV) 415\frac{4}{15}

Choose the correct answer from the options given below:

Answer options

Q12:

2025: 2 June Shift 1

Probability

Medium

Core

A problem in mathematics is given to three students whose chances of solving it are 1/2, 1/3, 1/4 respectively. The probability that the problem is solved is

Answer options

Q13:

2025: 2 June Shift 1

Probability

Medium

Core

In a college, 30% students fail in physics, 25% fail in Mathematics and 10% fail in both. One student is chosen at random. The probability that she fails in physics if she has failed in mathematics is

Answer options

Q14:

2025: 2 June Shift 1

Probability

Medium

Core

An urn contains 5 red and 5 black balls. A ball is drawn at random, its color is noted and is returned to the urn. Moreover, 2 additional balls of the same color are put in the urn and then a ball is drawn at random. The probability that the second drawn ball is red, is:

Answer options

Q15:

2025: 2 June Shift 1

Probability

Medium

Core

For any events A and B of a sample space S, which of the following statements are TRUE?

(A) P(S∣B)=1P(S | B) = 1

(B) P(A∩B)=P(A)+P(B)+P(A∪B)P(A \cap B) = P(A) + P(B) + P(A \cup B)

(C) P(Aˉ∣B)=1−P(A∣B)P(\bar{A} | B) = 1 - P(A | B)

(D) P(A∣B)=P(A∩B)P(B),P(B)≠0P(A | B) = \frac{P(A \cap B)}{P(B)}, P(B) \neq 0

Choose the correct answer from the options given below:

Answer options

Q16:

2025: 2 June Shift 1

Probability

Medium

Applied

The probability distribution of a random discrete variable is given

X-10123
P(X)0.1pp0.3qqrr

If it is known that P(X=1) is the mean of P(X=0) and P(X=2).

Then the value of r is :

Answer options

Q17:

2025: 2 June Shift 1

Probability

Medium

Applied

A pair of dice is thrown until the sum of numbers appeared is a perfect square or a non-perfect square sum appeared five times in succession. If random variable XX denotes the number of non perfect square sums appeared, then P(X>0)P(X > 0) is

Answer options

Q18:

2025: 2 June Shift 1

Probability

Hard

Applied

If the sum and difference of squares of mean and variance of a Binomial distribution is 225256\frac{225}{256} and 63256\frac{63}{256} respectively, the P(X≥2)P(X \geq 2) is:

Answer options

Q19:

2025: 2 June Shift 1

Probability

Medium

Applied

A and B are two independent events. The probability that both events A and B occur is 16\frac{1}{6} and the probability that neither of them occur is 13\frac{1}{3}. If P(A) = x, P(B) = y then the value of x+y is.

Answer options

Q20:

2025: 30 May Shift 2

Probability

Medium

Common

For a random variable x, probability distribution P(x) is given by P(x)=k6(3−x),x=0,1,2P(x) = \frac{k}{6}(3-x), x = 0, 1, 2, then

Match List-I with List-II

List-IList-II
(A) k is equal to(i) 12\frac{1}{2}
(B) P(x = 0)(ii) 1
(C) P(x < 2)(iii) 16\frac{1}{6}
(D) P(1 < x ≤ 2)(iv) 56\frac{5}{6}

Choose the correct answer from the options given below:

Answer options

Q21:

2025: 30 May Shift 2

Probability

Medium

Core

If A and B are two events such that P(A) ≠ 0 and P(B | A) = 1 then

Answer options

Q22:

2025: 30 May Shift 2

Probability

Medium

Core

Two numbers are selected without replacement at random, one at a time from the first six positive integers. Let x denotes the larger of the two numbers.

Match List-I with List-II

List-IList-II
(A) P(x = 2)(i) 415\frac{4}{15}
(B) P(x = 3)(ii) 115\frac{1}{15}
(C) P(x = 4)(iii) 215\frac{2}{15}
(D) P(x = 5)(iv) 15\frac{1}{5}

Choose the correct answer from the options given below:

Answer options

Q23:

2025: 30 May Shift 2

Probability

Medium

Core

Nitin has taken the subjects mathematics, physics and chemistry. The probability of him getting grade A in these subjects are respectively 0.2, 0.3 and 0.9. Getting grades in different subjects are regarded as independent events. The probability of getting A grade by him, either in mathematics or physics, is

Answer options

Q24:

2025: 30 May Shift 2

Probability

Medium

Core

If A and B are two events such that P(A)=12P(A) = \frac{1}{2}, P(B)=13P(B) = \frac{1}{3} and P(A∩B)=14P(A \cap B) = \frac{1}{4}, then which of the following statements are true?

(A) A and B are independent events

(B) P(A∣B)=34P(A | B) = \frac{3}{4}

(C) P(A′∣B′)=58P(A' | B') = \frac{5}{8}

(D) P(A′∣B)=14P(A' | B) = \frac{1}{4}

Choose the correct answer from the options given below:

Answer options

Q25:

2025: 30 May Shift 2

Probability

Medium

Applied

The probability of a shooter of hitting the target is 14\frac{1}{4}. The minimum number of fire needed so that the probability of hitting the target atleast once is greater than 716\frac{7}{16} is:

Answer options

Q26:

2025: 30 May Shift 2

Probability

Medium

Applied

Mean and variance of a binomial distribution are 6 and 2 respectively. The probability of 2 successes will be

Answer options

Q27:

2025: 30 May Shift 2

Probability

Easy

Applied

Probability distribution of random variable X is

X-2-1012
P(X)2/111/114/113/111/11

Then the value of E(X) is

Answer options

Q28:

2025: 30 May Shift 2

Probability

Medium

Applied

A random variable X has the following probability distribution:

X0123
P(X)0.20.10.30.4

The variance of X will be

Answer options

Q29:

2025: 30 May Shift 1

Probability

Medium

Common

The probability distribution of a random variable X is given by

X012
P(X)k2k3k

If k > 0, then P(0<X≤2)P(0 < X \leq 2) is equal to

Answer options

Q30:

2025: 30 May Shift 1

Probability

Medium

Core

The probability that A hits a target is 15\frac{1}{5} and the probability that B hits it is 23\frac{2}{3}. The probability that the target will be hit if both A and B shoot at it independently is:

Answer options